OpenStudy (anonymous):

solve (ab^2 + 10 + a) - (6ab^ - 2ab + 8)

OpenStudy (anonymous):

6ab^ something missing there

OpenStudy (anonymous):

6ab^2

OpenStudy (anonymous):

-5ab^2 + 2ab +2 + a

OpenStudy (anonymous):

solve (ab^2 + 10 + a) - (6ab^2 - 2ab + 8)

OpenStudy (anonymous):

you can use wolframalpha for such http://www.wolframalpha.com/input/?i=%28ab%5E2+%2B+10+%2B+a%29+-+%286ab%5E2+-+2ab+%2B+8%29

OpenStudy (mathteacher1729):

"solving" = find a particular value of a given variable. Your end result would be something like "x = 2" or "y = -pi" or something. "simplifying" = collecting like terms, cancelling common factors, etc. This problem is asking you to simplify, not solve. :)

OpenStudy (anonymous):

i need the actual math steps though

OpenStudy (anonymous):

then firstly remove brackets

OpenStudy (anonymous):

what

OpenStudy (anonymous):

-(a+b)=-a-b idk how to call this operation, but as you can see brackets are removed :P

OpenStudy (mertsj):

\[ab ^{2}+10 + a -(6ab ^{2}-2ab + 8)= ab ^{2} + 10 + a - 6ab ^{2} +2ab-8=-5ab ^{2}+2ab+a+2\]

OpenStudy (mertsj):

-5ab^2 +2ab+a+2

OpenStudy (mertsj):

Sorry, it got cut off.